# Shuchang Zhang, Hui Zhang, Hongxia Wang

### v4

@@ 4.2 A 2D toy example
We consider a $\frac{1}{\gamma}$-contractive operator $T_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from [41], for any $\mathbf{x}=(x_{1},x_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$,
@@
We compare the Picard iteration $\mathbf{x}^{k+1}=T_{\theta}\mathbf{x}^{k}$, PPP ($\lambda_{k}=0.1$), HPPP ($\mu_{k}=\frac{1}{k+1}$) and Restarted HPPP ($q=50$). Set $\mathcal{A}=\begin{pmatrix}-T_{\theta}&I\\ -I&I\end{pmatrix}$ and $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix}$ with $\tau=0.8,s=1.25$, and initial points $\mathbf{x}^{0}=\mathbf{y}^{0}=\mathbf{x}_{a}=\mathbf{y}_{a}=(1,0)^{\mathrm{T}}$, where $\mathbf{a}=\mathbf{u}^{0}$. Figure 4 shows that HPPP and Restarted HPPP indeed provide accelerated residual decay.
Figure 4: Trajectories, distance to solution, two residuals of iterates for the 2D toy example. Here $\gamma=1/0.95,\theta=15^{\circ},N=100$.
@@ 4.3 An infinite-dimensional example
We consider an infinite dimensional problem on $\ell^{2}$ from [7], the space of square summable sequences. For any $\mathbf{x}=(x_{0},x_{1},x_{2},\cdots)\in\ell^{2}$, the norm is $\left\|\mathbf{x}\right\|=\left(\sum_{k\in\mathbb{N}}\left|x_{i}\right|^{2}\right)^{\frac{1}{2}}$. We consider the following operator inclusion problem:
@@
where $\mathcal{S}(\mathbf{x})=(0,x_{0},x_{1},\cdots)$ is the right-shift operator and $\mathbf{b}=(1,-1,0,\cdots)$. Since $\mathcal{A}_{1},\mathcal{A}_{2}$ are nonexpansive, then $\mathcal{A}_{1},\mathcal{A}_{2}$ are maximal monotone [2, Example 20.29]. Additionally, by solving the above inclusion problem (22) we obtain $\mathbf{x}^{*}=(1,0,\cdots)$. Using $\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&I\end{pmatrix}$ and positive semi-definite preconditioner $\mathcal{M}=\begin{pmatrix}2I&-I\\ -I&\frac{1}{2}I\end{pmatrix}$, we compare the PPP iteration with $\lambda_{k}=0.5,1,1.8$. We set $N=10^{3},\tau=0.5,s=2$ and initialize $\mathbf{x}^{0}=(0.9,1,0,\cdots)$ for HPPP and PPP, and the anchor point is the same as the initial point, i.e., $\mathbf{a}=(\mathbf{x}^{0},\mathbf{x}^{0})$. For this example, we present visual results about the residual and the distance to the solution. As illustrated in Figure 5, HPPP outperforms PPP in terms of both $\left\|\mathbf{x}^{k}-\mathbf{x}^{*}\right\|$ and $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$.
@@
Figure 5: Comparison with PPP to solve an infinite dimensional problem in $\ell^{2}$ space.
@@ 4.4 Image deblurring
Firstly, we compared with CP [16], PPP [9], and HPPP for TV-regularized IR problems. For TV-regularized IR problems,
@@
We use a 2D Gaussian function with a standard deviation of 1.6 to convolve 10 test gray images, and finally obtain the degraded images with an additive white Gaussian noise with the noise level $\sigma=2.55$. Firstly, we compared three algorithms, CP, PPP, and the proposed HPPP. All the algorithms use the degraded images as initial points. We calculate the norm $K=\left\|\nabla\right\|=1.75$, and choose the total iteration $N=400$, balance coefficients $\lambda=2,\beta=5\times 10^{-4}$. Their parameters are given in Table 1. Both GraRED-P3 and GraRED-HP3 use DnCNN, other parameters are used below:
@@
Table 2 shows PSNR (dB) of restoration results on CP, PPP, HPPP, GraRED-P3 and GraRED-HP3. The performance of these methods is evaluated using PSNR measure. The best recovery results are highlighted in bold. From Table 2, GraRED-P3 and GraRED-HP3 are better than classic algorithms with explicit TV regularization, which demonstrates implicit regularizer is more powerful to regularize inverse imaging problems. We visualize the numerical comparison between GraRED-P3, GraRED-HP3, CP, PPP, and HPPP in Figure 6. To further compare the robustness of the initial points between the proposed HPPP, CP, and PPP with TV regularization. As shown in Figure 7, we plot their respective evolutions of PSNR values for iterations for Butterfly and Parrots with 10 different random initial points, here we choose the degraded image $\mathbf{y}$ as the anchor point. Once the anchor point is chosen, the proposed HPPP algorithm is robust to random perturbations of the initial points, which achieves stable recovery.
@@
Table 2: Deblurring results of gray images compared with CP, PPP, and HPPP about the Gauss blurring kernel.
@@
Secondly, we compared with RED and RED-PRO. Following [47, 20], a $9\times 9$ uniform point spread function (PSF) or a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images. We finally obtained the degraded images with the noise level $\sigma=\sqrt{2}$. The original RGB image is converted to the YCbCr image, all algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image. PSNR is measured on the luminance channel of the ground truth and the restored images. Table 3 shows the PSNR values of the restoration results on RED, RED-PRO, RRP, GraRED-P3, and GraRED-HP3. From the deblurring experiment Table 3 and Figure 9, GraRED-P3 and GraRED-HP3 achieve better performance than RED, RED-PRO, and RRP, which illustrates that KM or Halpern iteration used in PPP methods is effective. We further compare the differences among GraRED-P3, GraRED-HP3, and Restarted HPPP under six different settings with $\lambda=40,\tau=s=1,\lambda_{k}=0.2,\mu_{k}=\frac{1}{k+1},N=500,q=100$. In Table 5, we report the average PSNR values of 15 test images using six different blur kernels. When the anchor point and the initial point coincide, GraRED-HP3 achieves the best performance on some IR tasks. Moreover, by dynamically updating the anchor points, Restarted HPPP further boosts the performance of GraRED-HP3.
(a) Recovery peroformance
@@
Figure 8: Recovery peroformance and fixed-point residual comparison for the House image degraded by the Gaussian blurring kernel with the noise level $\sigma=5$. Here $D_{1}$ is used.
Additionally, we trained an FNE-DnCNN ($D_{1}$) using real spectral normalization [49], which is employed in RED and RED-PRO. We also compared with state-of-the-art methods using advanced denoisers such as DRUNet ($D_{2}$[65] and GS denoiser ($D_{3}$[32, 33, 59]. For GraRED-HP3, the anchor point is the same as the initial point. The default parameters are presented in Table 4. As shown in Table 6, Restarted HPPP with advanced denoisers achieves competitive or even superior performance compared to state-of-the-art methods. In Figure 8, we compare GraRED-P3 ($\lambda_{k}=0.1$) and GraRED-HP3 ($\mathbf{a}=\mathbf{u}^{0},\mu_{k}=\frac{1}{k+1}$). The results demonstrate that GraRED-HP3 achieves better recovery performance and exhibits a faster fixed-point convergence rate than GraRED-P3.
Table 3: Recovery results are obtained by different algorithms with TNRD [18]. The best two results are highlighted in red and blue colors, respectively.
@@
Table 5: Compared with GraRED-P3, GraRED-HP3, and Restarted HPPP with TNRD for different tasks.
@@
Table 6: Compared with state-of-the-art methods. Test gray images are degraded by the Gaussian kernel with the noise level $\sigma=2.55$. The best two results are highlighted in red and blue colors, respectively.
@@ 4.5 Image inpainting
We test 10 common images for evaluation. The first $\mathbf{M}$ is filled with a Bernoulli random mask whose each pixel is missing with probability $p=0.5$, i.e., $50\%$ of pixels are missed. The second $\mathbf{M}$ is a character mask where about $19\%$ of pixels are missed. All the algorithms start with the degraded image. For classic algorithms, we fix the balance parameter $\alpha=0.01$ and the total number $N=400$, and use the following other parameters:
@@
  • PPP: $\tau=s=1/\left\|K\right\|=0.57,\lambda_{k}=1.6$ or $\lambda_{k}=1.2$;
@@
Both GraRED-P3 and GraRED-HP3 use DnCNN [49], other parameters are used below:
@@
In Table 7 and Table 8, we compared the numerical performance of classic algorithms with TV regularization. In Figure 10 and 11, we compare visualization results of House degraded by the Bernoulli random mask and the character mask, the proposed algorithms achieve better visual performance than TV regularization.
@@ Appendix A Boundedness
Firstly, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5). To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce a important lemma from [30, Lemma 2.5].
@@
    $\{\mu_{k}\}$ converges to 0 in $[0,1]$, and $\sum_{k=0}^{\infty}\mu_{k}=+\infty$, or equivently $\prod_{k=0}^{\infty}(1-\mu_{k})=0$;
@@
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ be the sequence generated by (5). Then the following assertions are hold:
@@
Note that $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k}$, we have
@@ Appendix B $\mathcal{M}$-projection
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Then there exists a unique solution ${\mathbf{u}}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ which solves:
$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
@@
Since $l(\mathbf{u})=\frac{1}{2}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is a proper lower-semicontinuous differentiable convex function. Assume that $\mathbf{u}^{*}$ is the optimal solution of $\min_{\in\mathrm{Fix}(\mathcal{T})}l(\mathbf{u})$. Since $\mathrm{Fix}(\mathcal{T})=\mathrm{zer}\mathcal{A}$ is the convex set, thus $\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*})\in\mathrm{Fix}(\mathcal{T})$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$ and $t\in(0,1)$. Hence,
@@
Add above two inequalities, then we obtain $\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$. By (26), we have
@@
Assume $\mathbf{a}\in\mathcal{H}$, there exist an unique point $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ such that $\left\|\mathbf{u}^{*}-\mathbf{a}\right\|_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$, then $\mathbf{u}^{*}$ is called the $\mathcal{M}$-projection of $\mathbf{a}$ onto $\mathrm{Fix}(\mathcal{T})$, denoted by $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$.
@@
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Then the following assertions are equivalent:
@@
    $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$.

### v5

@@ 4.2 A 2D toy example
We consider a $\frac{1}{\gamma}$-contractive operator $T_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from [41] for any $\mathbf{x}=(x_{1},x_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$,
@@
We compare the Picard iteration $\mathbf{x}^{k+1}=T_{\theta}\mathbf{x}^{k}$, PPP ($\lambda_{k}=0.1$), HPPP ($\mu_{k}=\frac{1}{k+1}$), and restarted HPPP ($q=50$). Set $\mathcal{A}=\begin{pmatrix}-T_{\theta}&I\\ -I&I\end{pmatrix}$ and $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix}$ with $\tau=0.8,s=1.25$, and initial points $\mathbf{x}^{0}=\mathbf{y}^{0}=\mathbf{x}_{a}=\mathbf{y}_{a}=(1,0)^{\mathrm{T}}$, where $\mathbf{a}=\mathbf{u}^{0}$. Figure 4 shows that HPPP and restarted HPPP indeed provide accelerated residual decay.
Figure 4: Trajectories, distance to solution, and two residuals of iterates for the 2D toy example. Here $\gamma=1/0.95,\theta=15^{\circ},N=100$.
@@ 4.3 An infinite-dimensional example
We consider an infinite-dimensional problem on $\ell^{2}$ from [7], the space of square summable sequences. For any $\mathbf{x}=(x_{0},x_{1},x_{2},\ldots)\in\ell^{2}$, the norm is $\left\|\mathbf{x}\right\|=\left(\sum_{k\in\mathbb{N}}\left|x_{i}\right|^{2}\right)^{\frac{1}{2}}$. We consider the following operator inclusion problem:
@@
where $\mathcal{S}(\mathbf{x})=(0,x_{0},x_{1},\ldots)$ is the right-shift operator and $\mathbf{b}=(1,-1,0,\ldots)$. Since $\mathcal{A}_{1},\mathcal{A}_{2}$ are nonexpansive, $\mathcal{A}_{1},\mathcal{A}_{2}$ are maximal monotone [2, Example 20.29]. Additionally, by solving the above inclusion problem (22), we obtain $\mathbf{x}^{*}=(1,0,\ldots)$. Using $\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&I\end{pmatrix}$ and positive semidefinite preconditioner $\mathcal{M}=\begin{pmatrix}2I&-I\\ -I&\frac{1}{2}I\end{pmatrix}$, we compare the PPP iteration with $\lambda_{k}=0.5,1,1.8$. We set $N=10^{3},\tau=0.5,s=2$ and initialize $\mathbf{x}^{0}=(0.9,1,0,\ldots)$ for HPPP and PPP, and the anchor point is the same as the initial point, i.e., $\mathbf{a}=(\mathbf{x}^{0},\mathbf{x}^{0})$. For this example, we present visual results about the residual and the distance to the solution. As illustrated in Figure 5, HPPP outperforms PPP in terms of both $\left\|\mathbf{x}^{k}-\mathbf{x}^{*}\right\|$ and $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$.
@@
Figure 5: Comparison with PPP to solve an infinite-dimensional problem in $\ell^{2}$ space.
@@ 4.4 Image deblurring
First, we compared CP [16], PPP [9], and HPPP for TV-regularized IR problems. For TV-regularized IR problems,
@@
We use a 2D Gaussian function with a standard deviation of 1.6 to convolve 10 test grayscale images, and finally obtain the degraded images with an additive white Gaussian noise with the noise level $\sigma=2.55$. Firstly, we compared three algorithms, CP, PPP, and the proposed HPPP. All the algorithms use the degraded images as initial points. We calculate the norm $K=\left\|\nabla\right\|=1.75$, and choose the total iteration $N=400$, balance coefficients $\lambda=2,\beta=5\times 10^{-4}$. Their parameters are given in Table 1. Both GraRED-P3 and GraRED-HP3 use DnCNN; other parameters are used as follows:
@@
Table 2 shows PSNR (dB) of restoration results on CP, PPP, HPPP, GraRED-P3, and GraRED-HP3. The performance of these methods is evaluated using PSNR measure. The best recovery results are highlighted in bold. From Table 2, GraRED-P3 and GraRED-HP3 are better than classic algorithms with explicit TV regularization, which demonstrates that the implicit regularizer is more powerful for regularizing inverse imaging problems. We visualize the numerical comparison between GraRED-P3, GraRED-HP3, CP, PPP, and HPPP in Figure 6. We further compare the robustness of the initial points between the proposed HPPP, CP, and PPP with TV regularization. As shown in Figure 7, we plot their respective evolutions of PSNR values for iterations for Butterfly and Parrots with 10 different random initial points, here we choose the degraded image $\mathbf{y}$ as the anchor point. Once the anchor point is chosen, the proposed HPPP algorithm is robust to random perturbations of the initial points, which achieves stable recovery.
@@
Table 2: Deblurring results of grayscale images compared with CP, PPP, and HPPP about the Gauss blurring kernel.
@@
Second, we compared RED and RED-PRO. Following [47, 20], a $9\times 9$ uniform point spread function (PSF) and a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images. We finally obtained the degraded images with the noise level $\sigma=\sqrt{2}$. The original RGB image is converted into the YCbCr image, all algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image. PSNR is measured on the luminance channel of the ground truth and the restored images. Table 3 shows the PSNR values of the restoration results on RED, RED-PRO, RRP, GraRED-P3, and GraRED-HP3. From the deblurring experiment of Table 3 and Figure 9, GraRED-P3 and GraRED-HP3 achieve better performance than RED, RED-PRO, and RRP, which illustrates that KM or Halpern iteration used in PPP methods is effective. We further compare the differences among GraRED-P3, GraRED-HP3, and restarted HPPP under six different settings with $\lambda=40,\tau=s=1,\lambda_{k}=0.2,\mu_{k}=\frac{1}{k+1},N=500,q=100$. In Table 5, we report the average PSNR values of 15 test images using six different blur kernels. When the anchor point and the initial point coincide, GraRED-HP3 achieves the best performance on some IR tasks. Moreover, by dynamically updating the anchor points, restarted HPPP further boosts the performance of GraRED-HP3.
(a) Recovery performance
@@
Figure 8: Recovery performance and fixed-point residual comparison for the House image degraded by the Gaussian blurring kernel with the noise level $\sigma=5$. Here $D_{1}$ is used.
Additionally, we trained a FNE-DnCNN ($D_{1}$) using real spectral normalization [49], which is employed in RED and RED-PRO. We also compared our results with state-of-the-art methods using advanced denoisers, such as DRUNet ($D_{2}$[65] and GS ($D_{3}$[32, 33]. The PnP-DYS method [59] uses the Davis-Yin splitting (DYS) algorithm. For GraRED-HP3, the anchor point is the same as the initial point. The default parameters are presented in Table 4. As shown in Table 6, Restarted HPPP with advanced denoisers achieves competitive or even superior performance compared to state-of-the-art methods. In Figure 8, we compare GraRED-P3 ($\lambda_{k}=0.1$) and GraRED-HP3 ($\mathbf{a}=\mathbf{u}^{0},\mu_{k}=\frac{1}{k+1}$). The results demonstrate that GraRED-HP3 achieves better recovery performance and exhibits a faster fixed-point convergence rate than GraRED-P3.
Table 3: Recovery results are obtained by different algorithms with trainable nonlinear reaction-diffusion (TNRD) denoiser [18]. The best two results are highlighted in red and blue, respectively.
@@
Table 5: Comparison with GraRED-P3, GraRED-HP3, and restarted HPPP with TNRD for different tasks.
@@
Table 6: Comparison with state-of-the-art methods. Test grayscale images are degraded by the Gaussian kernel with the noise level $\sigma=2.55$. The best two results are highlighted in red and blue, respectively.
@@ 4.5 Image inpainting
We test 10 common images for evaluation. The first $\mathbf{M}$ is filled with a Bernoulli random mask, where each pixel is missing with probability $p=0.5$, i.e., $50\%$ of pixels are missed. The second $\mathbf{M}$ is a character mask where about $19\%$ of pixels are missed. All the algorithms start with the degraded image. For classic algorithms, we fix the balance parameter $\alpha=0.01$ and the total number $N=400$ and use the following additional parameters:
@@
  • PPP: $\tau=s=1/\left\|K\right\|=0.57,\lambda_{k}=1.6$, or $\lambda_{k}=1.2$;
@@
Both GraRED-P3 and GraRED-HP3 use DnCNN [49], other parameters are used:
@@
In Tables 7 and 8, we compared the numerical performance of classic algorithms with TV regularization. In Figures 10 and 11, we compare visualization results of House degraded by the Bernoulli random mask and the character mask; the proposed algorithms achieve better visual performance than TV regularization.
@@ Appendix A Boundedness
First, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5). To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce an important lemma from [30, Lemma 2.5].
@@
    $\{\mu_{k}\}$ converges to 0 in $[0,1]$, and $\sum_{k=0}^{\infty}\mu_{k}=+\infty$, or equivalently, $\prod_{k=0}^{\infty}(1-\mu_{k})=0$;
@@
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz. Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ be the sequence generated by (5). Then the following assertions are hold:
@@
Note that by $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k}$, we have
@@ Appendix B $\mathcal{M}$-projection
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz. Then there exists a unique solution ${\mathbf{u}}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ which solves:
$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0\quad\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
@@
Since $l(\mathbf{u})=\frac{1}{2}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is a proper lower-semicontinuous differentiable convex function, assume that $\mathbf{u}^{*}$ is the optimal solution of $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}l(\mathbf{u})$. Since $\mathrm{Fix}(\mathcal{T})=\mathrm{zer}\mathcal{A}$ is the convex set, thus $\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*})\in\mathrm{Fix}(\mathcal{T})$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$ and $t\in(0,1)$. Hence,
@@
Add the above two inequalities, we then obtain $\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$. By (26), we have
@@
Assume $\mathbf{a}\in\mathcal{H}$ and that there exists a unique point $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ such that $\left\|\mathbf{u}^{*}-\mathbf{a}\right\|_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$, then $\mathbf{u}^{*}$ is called the $\mathcal{M}$-projection of $\mathbf{a}$ onto $\mathrm{Fix}(\mathcal{T})$, denoted by $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$.
@@
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz. Then the following assertions are equivalent:
@@
    $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$.